if a 32 meter pole is broken into two parts at a ratio of 3 to 5, what is the measurement of the larger of the two parts
Answer by stanbon(75887)

if a 32 meter pole is broken into two parts at a ratio of 3 to 5, what is the measurement of the larger of the two parts
-----------------
Equation:
3x + 5x = 32
8x = 32
x = 4
------
larger = 5x = 5*4 = 20 meters
smaller = 3x = 3*4 = 12 meters
-----------
Cheers,
Stan H.
----------